how many liters of water must be added to 10L of 50% acid solution to produce 80% acid solution?
Found 2 solutions by ankor@dixie-net.com, drk:
Answer by ankor@dixie-net.com(22740)

how many liters of water must be added to 10L of 50% acid solution to produce 80% acid solution?
:
Impossible!
Water dilutes, you can only get less than 50% by adding water.

Answer by drk(1908)

This is a mixture problem. The thing to remember is that water is 0% everything. Here is the table:
liquid . . . . % . . . . Liter . . . . %Liter
acid . . . . . 50 . . . . 10 . . . . . 500
water . . . . 0 . . . . . L . . . . . 0
mixture . . . 80 . . . . 10 + L . . . 800 + 80L
The liters must be added to get the total amount of liquid.
Now,
500 + 0 = 800 + 80L
This is not possible. You can't add water and make it stronger. If it were 80% and you wanted a mixture of 50%, then we are good. Suppose that is what you meant. The new table looks like this:
liquid . . . . % . . . . Liter . . . . %Liter
acid . . . . . 80 . . . . 10 . . . . . 800
water . . . . 0 . . . . . L . . . . . 0
mixture . . . 50 . . . . 10 + L . . . 500 + 50L
Now,
800 + 0 = 500 + 50L
300 = 50L
L = 6. You would need 6 liters of water to dilute the mixture from 80 to 50%.